Take Control connection problems – resolved

1715hrs BST (GMT+1) / 1815hrs CET / 1215hrs EDT / 0915hrs PDT

The cause of the problem was tracked down to a SAN device. Although this had redundant storage processors, it was necessary to replace one of the storage processors in order for it to become operational again. We have an open ticket with our hardware vendor to determine why this was the case.

We do not believe this to be related to yesterday’s intermittent connectivity problem but cannot rule this and will raise with our hardware vendor.

We have been monitoring service availability for the last hour and full service would appear to be restored. Should you continue to experience any problems, please contact your local support team via Help Menu, Request Help in your Dashboard.

The GFI MAX Team.


This entry was posted in N-sight. Bookmark the permalink.